OpenStudy (minimato):

Find all solutions of the equation in the interval [0, 2pi). (sqrt3)csctheta+2 = 0. Write your answer in radians in terms of pi

OpenStudy (minimato):

\[\sqrt{3}\csc \theta+2=0\]

OpenStudy (phi):

I rarely use csc or sec, and so would first translate to 1/sin

OpenStudy (minimato):

Okay so \[\sqrt{3}(1/\sin \theta) + 2 = 0\]

OpenStudy (phi):

now add -2 to both sides, then divide both sides by sqr(3)

OpenStudy (phi):

you should get \[ \frac{1}{\sin \theta}= - \frac{2}{\sqrt{3}} \]

OpenStudy (phi):

now invert (i.e. "flip" ) both sides

OpenStudy (minimato):

Okay, got it! Sorry, was following along with you while writing it down on paper

OpenStudy (minimato):

Flip as in, turn upside down?

OpenStudy (phi):

yes, as in making the left side sin theta instead of 1/sin theta ditto for the right side.

OpenStudy (minimato):

\[\frac{ \sin \theta }{ 1 } = \frac{ -2 }{ \sqrt{3} }\]

OpenStudy (phi):

yes, but most people would just write sin theta on the left side

OpenStudy (phi):

oh, you have to flip the right side also

OpenStudy (minimato):

Oh, I forgot to flip the other side, sorry!

OpenStudy (minimato):

\[\sin \theta = -\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (phi):

usually, people memorize sin of certain angles: 0, 30, 45, 60 and 90 degrees this is one of those angles.

OpenStudy (phi):

*ignoring the sign, for the time being *

OpenStudy (minimato):

Is it 60? (ignoring the sign)

OpenStudy (phi):

yes. now we need to find the actual angle to get a general idea, we (maybe?) know that -sqr(3)/2 is about -.8 (roughly) if we plot a sin curve